OpenStudy (anonymous):

the sum of the squares of two consecutive integers is 41. Find the integers

OpenStudy (anonymous):

x+(x+1)=41 x+x=40 x=20 x+1=21 so 20 and 21

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

thats too high

OpenStudy (anonymous):

oh sheit i didnt read it sum of the sqrares so its probably x^2+(x+1)^2 answer is 4 and 5

OpenStudy (anonymous):

ok thx